Solutes do not typically exhibit ideal behavior in solution, resulting in non-integer van't Hoff factors.

Based on this information, which of the following solutes is most likely present in a 2.12 m aqueous solution with a boiling point of 103.87°C? (Kb for water is 0.512°C/m)

NaBr

C6H12O6

Na3PO4

Ca(ClO3)2
Answer #1

Change in boiling point from boiling point of water (100 °C), ∆Tb = 103.87 - 100 = 3.87 °C

Vant Hoff factor, i = ∆Tb/Kb×m

= 3.87/0.512×2.12

= 3.56

Since, vant Hoff factor is around 3.5 hence compound on dissociation give more than 3 ions. This is possible only for Na3PO4.

Answer is Na3PO4.